Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2016-04-05

Observing: Joel Johansson, Mark Magee, Ira Bar

Support Team: QUB

TAT: Lluís Galbany, Joe Lyman


Weather conditions


Start of night: cloudy, humidity ~38%, wind ~7 m/s

Middle of night: thin clouds, humidity ~36%, wind ~10m/s, seeing ~1.8"

End of night: thin clouds, humidity ~33%, wind ~8m/s, seeing ~1.5" - 2"



------------------------------------------------


Calibrations


Evening: PESSTO_bias, PESSTO_calib


------------------------------------------------


01:26 UT: Dome opened

01:37 UT: Follow up: PS15ae aborted due to bad seeing (>2")

01:51 UT: Follow up: SN2016aqf, gr13, 900s, slit 1.5"

02:16 UT: Follow up: SN2016O, gr13, 2x2700s, slit 1.5"

04:02 UT: Follow up: SN2016B, gr13, 1200s, slit 1.5"

04:30 UT: Spectroscopic standard: LTT3864

05:04 UT: Follow up: AT2016bln, gr13, 2700s, slit 1.5"

05:58 UT: Classification: AT2016blq/LSQ16ajh, gr13, 900s, slit 1.5"

06:22 UT: Classification: AT2016bls/LSQ16aki, gr13, 900s, slit 1.5"

06:44 UT: Follow up: SN2016ase, gr13, 1500s, slit 1.5"

07:19 UT: Spectroscopic standard: EG131

07:45 UT: Follow up: SN2015L, B x 120s, V x 60s

07:52 UT: Follow up: SN2015L, gr11, 2700s, slit 1.5"

08:42 UT: Follow up: SN2015L, gr16, 2700s, slit 1.5"

09:41 UT: Follow up: SN2015L, g x 120s, r x60s, i x60s

09:52 UT: Dome closed




Comments